# Learning Unified Reference Representation for Unsupervised Multi-class Anomaly Detection

### v1

@@ Abstract
In the field of multi-class anomaly detection, reconstruction-based methods derived from single-class anomaly detection face the well-known challenge of “learning shortcuts”, wherein the model fails to learn the patterns of normal samples as it should, opting instead for shortcuts such as identity mapping or artificial noise elimination. Consequently, the model becomes unable to reconstruct genuine anomalies as normal instances, resulting in a failure of anomaly detection. To counter this issue, we present a novel unified feature reconstruction-based anomaly detection framework termed RLR (Reconstruct features from a Learnable Reference representation). Unlike previous methods, RLR utilizes learnable reference representations to compel the model to learn normal feature patterns explicitly, thereby prevents the model from succumbing to the “learning shortcuts” issue. Additionally, RLR incorporates locality constraints into the learnable reference to facilitate more effective normal pattern capture and utilizes a masked learnable key attention mechanism to enhance robustness. Evaluation of RLR on the 15-category MVTec-AD dataset and the 12-category VisA dataset shows superior performance compared to state-of-the-art methods under the unified setting. The code of RLR will be publicly available.
@@
Unsupervised anomaly detection and localization strive to learn the patterns of normal samples from the training set then treat outliers as anomalies during inference, which is widely applied in industrial manufacturing [1, 34], medical image analysis [6], among other fields. However, in practical industrial anomaly detection scenarios, multi-class anomaly detection is not only more prevalent but also more valuable, as it only requires training one model for N classes, whereas single-class methods require training N models for N classes. Thus, this paper focuses on the promising and challenging multi-class anomaly detection.
@@ 2 Related Work
Transformer based Anomaly Detection. Recently, Transformer [24] has gained widespread attention for its ability to model long-range dependencies, leading to its utilization in reconstruction-based anomaly detection methods. InTra [16] employs a Transformer Encoder to reconstruct masked images. AnoViT [12] and VT-ADL [15] use Transformer Encoder to reconstruct features, which are further reconstructed into images using CNNs. ADTR [29] highlights Transformer Encoder-based methods suffer from “identity mapping” problem and proposes an Encoder-Decoder architecture for feature reconstruction. However, the outputs of these architectures are still primarily derived from the input mapping, leaving room for potential shortcuts in the learning process. This tendency becomes more pronounced, particularly in the case of multi-class anomaly detection, where the challenge of learning multi-class normal patterns increases, further reinforcing the inclination towards learning shortcuts, while our RLR addresses that.
@@ 3.2 Feature Reconstruction
To avoid “learning shortcut” issue described above, we replace self-attention in vanilla Transformer with our proposed parallel Masked Learnable Key Attention (MLKA) and Local Cross Attention (LCA). We combine the output of MLKA and LCA, where output of LCA multiplied by a hyperparameter greater than , to ensure model focuses more on LCA. Note that there is not residual shortcut between the input and the output of the attention module. Hence the new block is stacked by our proposed attention and Feed Forward Network (FFN) with norm & add similar to vanilla Transformer. For the input feature tokens at each feature level , after consecutive blocks, we obtain the recovered tokens , where is the number of patches of feature map. We formally define the calculation process of block as
@@
where is the output of the attention module, means the hidden layer dimension, represents LayerNorm, indicates the learnable reference tokens after projected to hidden feature space, and is the output of the block and the input of the block. Furthermore, we acquire from and from , which is defined as
@@
Specifically, we initialize a learnable reference feature representation for each feature level and define it as . Through training on the normal samples, the learnable reference eventually represents the normal feature pattern. We enhance the vanilla self-attention structure by applying the learnable normal reference feature tokens as the Key vectors. This allows us to eliminate the residual connection and prevent it from falling into rank collapse at the same time. Additionally, inspired by [28] we add a neighbor masked attention map to the original attention map between Query and Key vectors in order to make a token invisible to itself and its neighbors when calculating attention. The process of MLKA in block is formulated as
where parameters and are responsible for embedding and into Query, Key and Value vectors, and is the neighbor mask map that contains zero and negative infinity. In particular, the neighbor patches of the patch in are marked as negative infinity so that they will be zero after Softmax function to ignore patch’s neighbors. Since MLKA does not have residual connection and with neighbor mask, a token will not directly contain itself after passing through MLKA. Furthermore, due to the learnable key only consists of normal features, the abnormal tokens will receive a low similarity scores when calculating attention. As a result, the output of MLKA will not retain abnormal features. These advantages enable MLKA to recover abnormal tokens to their normal state and prevent the occurrence of “learning shortcuts”.
@@
where is the local aware mask similar to but the neighbor patches are marked as zero while others are negative infinity to focus on the neighbors. Since LCA forces the network to reconstruct the feature tokens from learnable reference representation, the reference eventually contains the normal features, thereby the recovered feature will get rid of abnormal features remarkably.
@@ 4.3 Main Results
Qualitative Results on MVTec-AD. To further demonstrate the superiority of our method, we visualized the reconstruction and detection results of the MVTec-AD [1] dataset, as shown in Figure 3. It can be observed that UniAD fails to detect some anomalies as these anomaly features are not recovered to normal features. In contrast, our method successfully reconstructs normal features. This indicates that UniAD has not fully learned the normal patterns for these classes but instead learned shortcuts that prevent the reconstruction of anomalies as normal features. In contrast, our method recovers these samples to normal instances, demonstrates the effectiveness of our RLR in successfully addressing the issue of models learning shortcuts.
Qualitative Results on VisA. We also conducted a qualitative analysis on the VisA [34] dataset. Figure 4 presents the visual detection results of UniAD [28] and our RLR on the VisA dataset. It is important to note that we were unable to reproduce the results of OmniAL [33] due to the unavailability of its source code, hence its results are not shown. Similar to the qualitative analysis conducted on the MVTec-AD [1] dataset, our method successfully detects samples that UniAD fails to detect, thus demonstrating the effectiveness of our approach.

### v2

@@ Abstract
In the field of multi-class anomaly detection, reconstruction-based methods derived from single-class anomaly detection face the well-known challenge of “learning shortcuts”, wherein the model fails to learn the patterns of normal samples as it should, opting instead for shortcuts such as identity mapping or artificial noise elimination. Consequently, the model becomes unable to reconstruct genuine anomalies as normal instances, resulting in a failure of anomaly detection. To counter this issue, we present a novel unified feature reconstruction-based anomaly detection framework termed RLR (Reconstruct features from a Learnable Reference representation). Unlike previous methods, RLR utilizes learnable reference representations to compel the model to learn normal feature patterns explicitly, thereby prevents the model from succumbing to the “learning shortcuts” issue. Additionally, RLR incorporates locality constraints into the learnable reference to facilitate more effective normal pattern capture and utilizes a masked learnable key attention mechanism to enhance robustness. Evaluation of RLR on the 15-category MVTec-AD dataset and the 12-category VisA dataset shows superior performance compared to state-of-the-art methods under the unified setting. Code is available at RLR.
@@
Unsupervised anomaly detection and localization strive to learn the patterns of normal samples from the training set then treat outliers as anomalies during inference, which is widely applied in industrial manufacturing [1, 37, 25, 33], medical image analysis [6], among other fields. However, in practical industrial anomaly detection scenarios, multi-class anomaly detection is not only more prevalent but also more valuable, as it only requires training one model for N classes, whereas single-class methods require training N models for N classes. Thus, this paper focuses on the promising and challenging multi-class anomaly detection.
@@ 2 Related Work
Transformer based Anomaly Detection. Recently, Transformer [24, 34] has gained widespread attention for its ability to model long-range dependencies, leading to its utilization in reconstruction-based anomaly detection methods. InTra [16] employs a Transformer Encoder to reconstruct masked images. AnoViT [12] and VT-ADL [15] use Transformer Encoder to reconstruct features, which are further reconstructed into images using CNNs. ADTR [30] highlights Transformer Encoder-based methods suffer from “identity mapping” problem and proposes an Encoder-Decoder architecture for feature reconstruction. However, the outputs of these architectures are still primarily derived from the input mapping, leaving room for potential shortcuts in the learning process. This tendency becomes more pronounced, particularly in the case of multi-class anomaly detection, where the challenge of learning multi-class normal patterns increases, further reinforcing the inclination towards learning shortcuts, while our RLR addresses that.
@@ 3.2 Feature Reconstruction
To avoid “learning shortcut” issue described above, we replace self-attention in vanilla Transformer with our proposed parallel Masked Learnable Key Attention (MLKA) and Local Cross Attention (LCA). We combine the output of MLKA and LCA, where output of LCA multiplied by a hyperparameter greater than , to ensure model focuses more on LCA. Note that there is not residual shortcut between the input and the output of the attention module. Hence the new block is stacked by our proposed attention and Feed Forward Network (FFN) with norm & add similar to vanilla Transformer. For the input feature tokens at each feature level , after consecutive blocks, we obtain the recovered tokens , where is the number of patches of feature map, i.e., . We formally define the calculation process of block as
@@
where is the output of the attention module, means the hidden layer dimension, represents LayerNorm, indicates the learnable reference tokens after projected to hidden feature space, and is the output of the block and the input of the block. We use a single for all blocks, so it is independent of . Furthermore, we acquire from and from , which is defined as
@@
Specifically, we randomly initialize a learnable reference feature representation for each feature level and define it as , which has the same size as because their patch positions need to correspond. Then we use a fully connected layer to project to the hidden feature space into , and a single is used for all K layers. Through training on the normal samples, the learnable reference eventually represents the normal feature pattern. We enhance the vanilla self-attention structure by applying the learnable normal reference feature tokens as the Key vectors. This allows us to eliminate the residual connection and prevent it from falling into rank collapse at the same time. Additionally, inspired by [29] we add a neighbor masked attention map to the original attention map between Query and Key vectors in order to make a token invisible to itself and its neighbors when calculating attention. The process of MLKA in block is formulated as
where parameters and are responsible for embedding and into Query, Key and Value vectors, and is the neighbor mask map that contains zero and negative infinity. In particular, the neighbor patches of the patch in are marked as negative infinity so that they will be zero after Softmax function, meaning the attention weights of neighbor patches to the patch in are zero, aiming to ignore patchi’s neighbors. The size of the neighbor window for each feature level is a hyperparameter, and we provide specific settings in the experimental section. Since MLKA does not have residual connection and with neighbor mask, a token will not directly contain itself after passing through MLKA. Furthermore, due to the learnable key only consists of normal features, the abnormal tokens will receive a low similarity scores when calculating attention. As a result, the output of MLKA will not retain abnormal features. These advantages enable MLKA to recover abnormal tokens to their normal state and prevent the occurrence of “learning shortcuts”.
@@
where is the local aware mask similar to but the neighbor patches are marked as zero while others are negative infinity, meaning the attention weights of non-neighbor patches to the patch in are zero, thereby focusing attention only on the neighbor patches. Since LCA forces the network to reconstruct the feature tokens from learnable reference representation, the reference eventually contains the normal features, thereby the recovered feature will get rid of abnormal features remarkably.
@@ 4.3 Main Results
Qualitative Results on MVTec-AD. To further demonstrate the superiority of our method, we visualized the reconstruction and detection results of the MVTec-AD [1] dataset, as shown in Figure 3. It can be observed that UniAD fails to detect some anomalies as these anomaly features are not recovered to normal features. In contrast, our method successfully reconstructs normal features. This indicates that UniAD has not fully learned the normal patterns for these classes but instead learned shortcuts that prevent the reconstruction of anomalies as normal features. In contrast, our method recovers these samples to normal instances, demonstrates the effectiveness of our RLR in successfully addressing the issue of models learning shortcuts. We also conducted qualitative analysis on the VisA [37] dataset. Due to space limitations, please refer to the supplementary materials for details.
@@ 5 Conclusion
Acknowledgements
This work was supported by National Natural Science Foundation of China, No.62171139.
Supplementary Material
1 Overview
2 Fusion of MLKA and LCA
In order to integrate the outputs of MLKA and LCA modules, we explored alternative fusion methods. In addition to employing the hyperparameter to perform a weighted summation of their outputs, we conducted experiments by introducing learnable parameters to facilitate the weighted summation of the two modules. This modification involved transforming Equation (4) in the main text into
where and are the parameters learned by the model. The experimental results are presented in Table 1, which demonstrate that the use of learnable weights increases the likelihood of the model learning shortcuts, potentially focusing more on MLKA rather than LCA. This suggests that by using the hyperparameter , we can encourage the model to pay more attention to LCA, thereby compelling it to learn normal patterns and achieve better detection performance.
| Fusion | w/ | w/ |
| Metrics | 98.1 / 98.2 | 98.6 / 98.5 |
3 Qualitative Results on VisA.
To further validate the superiority of our approach, we also conducted a qualitative analysis on the VisA [37] dataset. Figure 1 presents the visual detection results of UniAD [29] and our RLR on the VisA dataset. It is important to note that we were unable to reproduce the results of OmniAL [36] due to the unavailability of its source code, hence its results are not shown. Similar to the qualitative analysis conducted on the MVTec-AD [1] dataset, our method successfully detects samples that UniAD fails to detect, thus demonstrating the effectiveness of our approach.
4 Multiple Learnable References
We conducted experimental analysis on the quantity of learnable references. Specifically, we performed calculations of attention with multiple references in the LCA and summed the outputs, which transformed in Equation (4) in the main text into , where is the number of learnable references. Additionally, we imposed constraints on the diversity of references in the loss function. The experimental results are presented in Table 2, which indicates that this approach does not lead to performance improvement; instead, it incurs higher computational and storage resource consumption. Therefore, this paper only utilizes a single learnable reference representation ().
| 1 | 2 | 3 | |
| Metrics | 98.6 / 98.5 | 98.6 / 98.5 | 98.6 / 98.4 |
5 The Failure Analysis
We conducted a failure analysis on classes with low performance metrics on the MVTec-AD and VisA datasets. Specifically, in MVTec-AD, the Toothbrush class exhibits lower metrics due to the recognition of noise in the background as anomalies, resulting in some normal samples being misclassified as abnormal samples. This phenomenon is more likely to occur in feature reconstruction methods, as evidenced by similarly low metrics for this class in UniAD. Additionally, in the VisA dataset, the Capsules class exhibits lower metrics owing to the presence of random discolorations. Reconstruction models tend to confuse these discolorations with subtle anomalies, leading to poor performance of RLR and UniAD on this class.